# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v2

@@ Abstract
I Introduction
@@
Managing energy assets in a grid system is a complex task that requires decision-making under uncertainty. The dynamics and stochasticity of the environment create a constantly evolving and inherently complex system. This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization.
To effectively manage this, it is useful to combine the forecast of the problem environment with the optimization of the decision-making process. This combination of forecasting and decision-making is often referred to as the predict, then optimize problem, or an online optimization with predictions. The online optimization problems with predictions are commonly seen in energy applications such as electric vehicle charging [1] [2], battery scheduling [3][4] and energy and flexibility dispatch problems [5] [6]. In number of these applications, the decision-maker is faced with switching costs, which are the costs associated with updating the operational plan. For instance, transitioning from $x_{t-1}$ to $x_{t}$ typically incurs a switching cost or ramp cost, $d(x_{t}-x_{t-1})$. Consequently, the stage cost becomes time-coupled and is defined as: $C_{t}(x_{t-1},x_{t}):=f_{t}(x_{t})+d(x_{t}-x_{t-1})$.
In the energy sector, the switching costs can arise from various factors, such as the physical limitations of ramping up or down energy production [7], managing server load balancing [8] and payment of network fees associated with energy trading [9]. While there has been a significant amount of research on online optimization problems with predictions, there is a gap in the literature on the impact of switching costs on the interaction between forecast model and the optimization algorithm. We address this gap by studying two aspects of the combined system: time-coupling and forecast stability.
Switching costs are a direct consequence of time-coupling, when the action $x_{t}$ at time $t$ influences the future actions $x_{t+1}$ and system states $S_{t+1}$, as defined by Lin et al. [8]. Errors in forecasts act as ’switching incentive‘ compelling policy revisions using more accurate predictions aligned with up-to-date information and the actual state of the system. This perspective suggests a balancing act between forecast accuracy and switching costs. This paper explores how this trade-off influences the system’s design and operation, suggesting a nuanced view of forecasting’s role in optimization.
In principle, forecasts used in optimization can be updated every time new information becomes available, often span multiple steps ahead and are revised periodically. Transitioning from the concept of time-coupling, we speculate that switching costs not only underscore the importance of forecast accuracy but also elevate the value of forecast stability. In this context, a stable forecast with minimal variation despite periodic updates, minimizes the need for frequent policy changes, effectively reducing switching costs. Consequently, this research aims to answer the following questions:
@@
  • And how can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
I-A Related work
@@
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [13], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model.
@@
We are particularly interested in the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm, also referred to as model predictive control (MPC). When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The research conducted by Chen et al. [15] sheds light on the role of noisy prediction in the design of the OCO algorithm. Those authors propose a model that simulates realistic prediction errors. Using this model, they show analytically that an alternative Averaging Fixed Horizon Control (AFHC) algorithm can achieve sublinear regret and a constant competitive ratio, which indicate a desirable level of optimality in reference to the theoretical best policy. Figure 1 shows an overview of the RHC and AFHC algorithm designs, along with other algorithms discussed further. Those authors state that other algorithms can reach better performance if they can use the predictions more efficiently. A later work, Chen et al. [16], introduces a class of policies known as Committed Horizon Control (CHC), which extends the principles of Receding Horizon Control (RHC) and Adaptive Forward Horizon Control (AFHC) through the integration of a commitment concept. This concept involves specifying a predetermined number of steps that the algorithm executes. The analysis of the generalized problem reveals a trade-off set by the varying levels of switching costs and quality of the predictions. The general observation is that the level of commitment is problem-specific and depends on the nature of the switching costs and the quality of the predictions. Therefore, the results show that the commitment level should be tuned for maximizing the performance of the policy.
(a) RHC
(b) FHC
(c) AFHC
(d) CHC
@@
Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control, AFHC: Averaging Fixed Horizon Control, CHC: Committed Horizon Control.
Current work sets to explore the performance of commitment-based policies in the context of energy management systems. We are particularly interested in analyzing the forecast and planning stability, as a potential factor that influences the performance of the policy. A comprehensive study of forecast stability is made by Godahewa et al. [17]. Those authors categorize the stability into two types: vertical and horizontal. Vertical stability refers to the variation period-to-period variability between forecasts given for the same time period but issued from different origins. Horizontal stability describes the variation between predictions made at one instance within a prediction window. The lack of planning stability has long been recognized an important issue within research on supply chain management. The vertical stability in order forecasts is often referred to as the system nervousness. It is a well-known phenomenon that can lead to inefficient production and inventory management. The extra capacity implications are also referred to as the bullwhip effect [18]. The commonly used methods to obtain a stable plan include increased commitment, referred as freezing intervals [19] [20], using smoothing techniques [21], and incorporating changing costs into the objective function [22] [23].
Unlike supply chain management, energy management systems are more dynamic and require more frequent updates of forecasts and policies. Switching costs are also observed in energy, such as the ramping costs of generating units [7] [24] and toggling server load balancing [8] [25]. Furthermore, the literature is scarcer when considering the operational aspects of a combined forecasting and optimization problem. Guo et al. [26] review the historical evolution and state-of-the-art for the optimization methods for energy storage dispatch. They conclude that current research indicates that RHC performance relies strongly on the quality of predictions, and therefore, it may not be robust to uncertainty. Christianson et al. [27] analyze the problem of resource procurement and online economic dispatch for power system operators. They present a customized FHC algorithm — Feasible Fixed Horizon Control (FFHC) — which incorporates a regularization term to solve the multi-interval optimization problem while ensuring feasibility and robustness amidst the complexities introduced by ramping costs. A study by La Tona et al. [28] explores how the frequency of forecast revisions affects the Energy Management System’s (EMS) performance in online replanning tasks, focusing on battery control to align actual and declared consumption. Their findings suggest that more frequent forecasts enhance EMS performance, with diminishing returns at intervals of 60 minutes and shorter. The current work examines applications in energy storage optimization and the relationship between forecast and optimal commitment level.
I-B Contents and Contributions
Through a systematic review of the literature, it becomes evident that the effects of predictions on online optimization problems with switching costs are not well understood. This present research aims to address this gap by studying the impact of forecast stability and accuracy on the performance of the policy in the context of energy management systems. The principle contributions are as follows:
@@ I-B Contents and Contributions
  • We investigate the effect of forecast accuracy and stability on the downstream performance of the optimization policy. We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
  • An empirical evaluation of the stability of deterministic and probabilistic scenario sets is conducted. We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy. We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
@@
II Methodology
II-A Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [29] and based on a real-world dataset from a grid-interactive community. More generally, we consider an online optimization with switching costs. The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Online Appendix A [^1] Appendix A. [Online] Available: https://arxiv.org/abs/2407.03368. The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [30], we use the following formulation:
$X_{t}^{MPC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
@@ II-A Problem statement
  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$. This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
@@
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Online Appendix A  [^2] Appendix A. [Online] Available: https://arxiv.org/abs/2407.03368. Furthermore, In the Online Appendix C  [^3] Appendix C. [Online] Available: https://arxiv.org/abs/2407.03368, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
II-B Optimization
@@
Similar to the optimization, the forecast is also run in a rolling origin fashion. The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps. In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps. Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. The resulting decision matrix is shown in Table I.
TABLE I: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
@@ II-C Forecast
II-C1 Rolling Origin Window
@@
II-C2 Stability
As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast. The stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed by Godahewa et al. [17], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 3. We define the vertical stability as the difference between the predictions produced at different origins for the same time period, marked with the overlap area in Figure 2. If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal. Horizontal stability characterizes the difference between predictions generated at the same origin. The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
@@ II-C2 Stability
Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions
Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC). The formulations of the stability metrics vary for vertical and horizontal types. In line with the definition of sMAPC in [34], we define the absolute scale stability metric, Mean Absolute Change (MAC). The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
@@
II-C3 Stability of Probabilistic Forecasts
The Earth Mover’s Distance (EMD), or Wasserstein distance, quantifies the minimum ”work” required to transform one distribution into another, making it ideal for assessing the stability of scenario sets in stochastic optimization. EMD is chosen for its symmetric nature, adherence to the triangle inequality, and ability to compare distributions on different probability spaces, making it effective even when forecasts differ significantly. It also accounts for the metric space dimension, akin to the mean absolute change (MAC) used for point forecasts. Other metrics like Kullback-Leibler (KL) divergence and Jensen-Shannon distance measure statistical distance, but EMD is preferred for scenario sample sets. In a 1-dimensional case with empirical sets $P$ and $Q$, EMD is defined as:
$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|P_{(i)}-Q_{(i)}\right\|$ (6)
where $P_{(i)}$ and $Q_{(i)}$ are the sorted samples of $P$ and $Q$ respectively. This metric measures the dissimilarity between two sets of scenarios. To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability:
$EMD_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (7)
$EMD_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (8)
where $\hat{y}_{t+i\mid t,j}$ is the $j$-th scenario of the forecast for time step $t+i$ generated at time $t$, and $N$ is the number of scenarios.
II-C4 Accuracy of Determinstic Forecasts
For deterministic forecasts, the MAE is utilized to quantify the forecast accuracy. It calculates the average absolute difference between each forecasted value and the corresponding actual value. Mathematically, it is expressed as:

### v3

@@ Abstract
{highlights}
The paper evaluates the impact of forecast stability and accuracy on the downstream performance of energy management systems under switching costs.
A novel metric, Scenario Distribution Change (SDC), is proposed to assess the stability of probabilistic scenario sets, extending the Wasserstein distance.
Forecast accuracy and stability are shown to influence the optimal policy commitment level. There is a theoretical trade-off between switching costs and forecast errors.
Empirical analysis is conducted using the CityLearn 2022 competition case study, highlighting the trade-offs between forecast updates, stability, and operational outcomes. The longer than minimal commitment periods are found to improve system performance in case of stochastic optimization with probabilistic scenario sets.
1 Introduction
@@
Managing energy assets within a grid system presents a challenging task characterized by decision-making under uncertainty. The inherent dynamics and stochasticity of the environment make this a complex system, constantly evolving and demanding decisions to be made with incomplete knowledge of the future over limited time windows, a scenario typically addressed through online optimization.
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information. Striking a balance between forecast accuracy and the impact of switching costs is therefore crucial. This paper explores how this balance shapes system design and informs operational strategies.
Building on this interplay between forecast accuracy and switching costs, we propose that the stability of forecasts plays an equally critical role. Since forecasts in optimization are typically updated with each new data acquisition—covering multiple future steps and undergoing regular revisions—stability becomes essential. A forecast that remains consistent across updates reduces the need for frequent strategic adjustments, thereby mitigating the impact of switching costs. Consequently, this research aims to answer the following questions:
  • 1.
@@
  • 3.
    How can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
1.1 Related work
@@
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [13], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
@@
We are particularly interested in the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
(a) FHC
(b) RHC
@@
Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control
The research conducted by Chen et al. [15] highlights the challenges posed by noisy predictions in designing Online Convex Optimization (OCO) algorithms. Several solutions, such as Averaging Fixed Horizon Control (AFHC) and its generalized counterpart, Committed Horizon Control (CHC) [16], propose aggregation strategies over multiple revisions of plans to achieve sublinear regret and improve robustness to prediction errors. Similarly, the Feasible Fixed Horizon Control (FFHC) algorithm [17] incorporates a regularization term to handle multi-interval optimization problems while ensuring feasibility and robustness against ramping costs.
While these approaches demonstrate the benefits of aggregating policies over prediction updates, the current work shifts focus to understanding the effects of commitment on forecast stability and downstream optimization in the presence of switching costs. Unlike these aggregation-based methods, this study seeks to explore how the commitment level impacts system performance, with a particular emphasis on the relationship between forecast stability and switching costs.
Forecast stability has been recognized as a critical factor in planning. Building on the comprehensive framework by Godahewa et al. [18], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems. Unlike supply chain management, where freezing intervals [19, 20] and smoothing techniques [21] have been widely used to address system nervousness and the bullwhip effect [22], energy systems require more dynamic and frequent forecast updates.
Nevertheless, switching costs are observed in energy management systems, such as ramping costs for generating units [7, 23] or toggling server load balancing [8, 24]. However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs. By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
1.2 Contents and Contributions
The principle contributions are as follows:
  • 1.
@@ 1.2 Contents and Contributions
  • 2.
    The effect of forecast accuracy and stability on the downstream performance of the optimization policy is analyzed. We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
  • 3.
    We conduct an empirical evaluation of the stability of point forecasts and probabilistic scenario sets, using deterministic and stochastic optimization approaches, respectively. We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy. We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
  • 4.
@@
2 Methodology
2.1 Problem statement
Figure 2: Grid diagram illustrating the coordinated battery scheduling problem.
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [25] and based on a real-world dataset from a grid-interactive community. The setup of the problem is shown in the Figure 2. More generally, we consider an online optimization with switching costs. The problem is solved with the fixed horizon control (FHC), the variation of model predictive control (MPC) with a fixed commitment horizon. The full details of the objective score and relevant constraints are provided in A.1. The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [26], we use the following formulation:
$X_{t}^{FHC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
@@ 2.1 Problem statement
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$. This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
@@
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in A.2. Furthermore, In A.3, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. Following the line of proof in Chen et al. [16], we derive the upper bound on the competitive difference between the FHC algorithm and the optimal policy. The analysis is conducted for the deterministic case, where the forecast is a point forecast, then extended to the probabilistic case, where the forecast is represented as a set of scenarios. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
2.2 Optimization
@@
Similar to the optimization, the forecast is also run in a rolling origin fashion. The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps. In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps. Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. The resulting decision matrix is shown in Table 1.
Table 1: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
@@ 2.3 Forecast
2.3.1 Rolling Origin Window
@@
2.3.2 Stability
As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast. The stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed in [18], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 4.
@@ 2.3.2 Stability
Figure 4: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions. The vertical stability is the difference between the predictions produced at different origins for the same time period. If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal. Horizontal stability characterizes the difference between predictions generated at the same origin. The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
Similarly to the evaluation in [30], a stability metric is measured for point forecasts using the mean absolute change (MAC). The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the change of predictions within the horizon window generated at the same origin — ${MAC}_{H}$
@@
2.3.3 Stability of Probabilistic Forecasts
In stochastic optimization, scenario stability measures the consistency of probabilistic forecast updates over time. We propose the Scenario Distribution Change (SDC) metric, adapting the Wasserstein Distance, also referred to as Earth Mover Distance, to assess the stability of scenario sets. SDC quantifies the average change between successive scenario updates, providing a clear measure of stability for probabilistic forecasts in dynamic systems.
The Wasserstein distance formulation is symmetric, adheres to the triangle inequality, and effectively compares probabilistic scenario sets. Unlike metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon distance, SDC holds for non-overlapping distributions, making it particularly relevant for applications requiring forecast stability. By incorporating the dimension of the metric space, SDC is conceptually analogous to the Mean Absolute Change (MAC) used for point forecasts.
To assess stability across time, SDC evaluates both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates). For $N$ number of scenarios, these are defined as:
Horizontal Stability (SDCH):
$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (6)
Vertical Stability (SDCV):
$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (7)
2.3.4 Accuracy of Determinstic Forecasts
For point forecasts, the MAE is utilized to quantify the forecast accuracy. It calculates the average absolute difference between each forecasted value and the corresponding actual value. Mathematically, it is expressed as: